### v3

# The Critical LYZ equation in Kähler Geometry

Abstract.
We establish the existence of smooth solutions for the LYZ equation at the critical phase $\theta=(n-2)\frac{\pi}{2}$, thereby solving the critical case of a problem posed by Collins-Jacob-Yau [7] and Li [26] concerning the solvability for phase $\theta\leq(n-2)\frac{\pi}{2}$. As applications, we solve the 3D Hessian equation $\sigma_{2}=1$ and the 4D Hessian quotient equation $\sigma_{3}=\sigma_{1}$ under weaker assumptions than previously required [36, 38].
Keywords: critical LYZ equation; subsolution; uniform estimates; blow-up argument
footnotetext: 2020 Mathematics Subject Classification. Primary 32W20; Secondary 53C55.
1. Introduction
In Kähler geometry, the LYZ (Leung-Yau-Zaslow) equation, also known as the dHYM (deformed Hermitian Yang-Mills) equation, is a fully nonlinear elliptic PDE that arises from mirror symmetry [25, 31]. This equation is of current interest in both mathematics and physics.
Let $(M,\omega)$ be a compact Kähler manifold and $\chi$ a closed real $(1,1)$-form. Suppose that $\mathcal{Z}=\int_{M}(\chi+\sqrt{-1}\omega)^{n}\neq 0$ and let $\hat{\theta}$ be the principal argument of $\mathcal{Z}$. The LYZ equation seeks a smooth function $u:M\rightarrow\mathbb{R}$ such that
$\mathrm{Im}\big(e^{-\sqrt{-1}\hat{\theta}}(\chi_{u}+\sqrt{-1}\omega)^{n}\big)=0.$ (1)
Let $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ and define the special Lagrangian operator as
$\theta_{\omega}(\chi_{u}):=\sum\limits_{i=1}^{n}\arctan\lambda_{i}.$
Then equation (1) can be rewritten in the form
$\theta_{\omega}(\chi_{u})=\theta$ (2)
where $\theta=n\frac{\pi}{2}-\hat{\theta}$. We refer to (2) as the supercritical LYZ equation if $\theta\in\big((n-2)\frac{\pi}{2},n\frac{\pi}{2}\big)$, and as the critical LYZ equation if $\theta=(n-2)\frac{\pi}{2}$, i.e., $\hat{\theta}=\pi$.
A key requirement for solving the LYZ equation is the existence of a subsolution: A smooth function $\underline{u}$ satisfying
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>\theta-\frac{\pi}{2}.$ (3)
This notion of subsolution is equivalent to that introduced by Guan [13] and Székelyhidi [38].
1.1. Previous Results
The 2-dimensional case of the LYZ equation was solved by Jacob-Yau [23], who also obtained partial results in higher dimensions. Collins-Jacob-Yau [7] then solved the supercritical case under the assumption of a subsolution $\underline{u}$ and an additional condition $\theta_{\omega}(\chi_{\underline{u}})>(n-2)\frac{\pi}{2}$. The authors [12] introduced a geometric flow to solve the equation under the same conditions. The extra condition was subsequently removed by Pingali [34] for $n=3$, by Lin [28] for $n=3,4$, and finally by Lin [27] in full generality. An alternative proof, based on Chen’s significant result [3], was given by Sun [37]. For further recent developments, we refer to [2, 4, 5, 6, 8, 9, 14, 15, 16, 19, 20, 21, 22, 24, 29, 39].
1.2. Our main results
Collins-Jacob-Yau [7] suggested studying the LYZ equation when $\theta\leq(n-2)\frac{\pi}{2}$. This was later explicitly posed as a question by Li [26], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$, the bounded derived category of coherent sheaves on $M$.
In this paper, we solve the critical case; that is, we establish the existence of smooth solutions for the critical LYZ equation:
$\theta_{\omega}(\chi_{u})=(n-2)\frac{\pi}{2}.$ (4)
In this case, the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{n}$ is $\pi$, and the subsolution condition becomes
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>{(n-3)\frac{\pi}{2}}.$ (5)
Theorem 1.1.
Let $(M,\omega)$ be a compact Kähler manifold and $\chi$ a closed real (1,1)-form such that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{n}$ is $\pi$. If there exists a smooth function $\underline{u}$ satisfying (5), then the critical LYZ equation (4) admits a unique smooth solution u with $\sup_{M}u=0$.
Remark 1.1.
For the special Lagrangian equation in $\mathbb{R}^{n}$, interior Hessian estimates for critical and supercritical cases were proved by Warren-Yuan [42] and Wang-Yuan [41]. For $\theta<(n-2)\frac{\pi}{2}$, $C^{1,\alpha}$ solutions were constructed by Nadirashvili-Vlădut [33] and Wang-Yuan [40]. Very recently, Mooney-Savin [32] obtained a Lipschitz solution that is not $C^{1}$.
To solve the critical LYZ equation (4), we consider the family of equations:
$\theta_{\omega}(\chi_{u^{t}}+t\omega)=\theta(t).$ (6)
We verify that there exists a uniform constant $C>0$ such that
$\theta(t)\in((n-2)\frac{\pi}{2}+Ct,(n-2)\frac{\pi}{2}+2Ct)$
for small $t>0$ (see Lemma 2.3 in Section 2). Hence equations (6) are the supercritical LYZ equations. Then we prove the subsolution $\underline{u}$ of the critical LYZ equation (4) is also a subsolution of equations (6) for small $t>0$. Thus we can apply Lin’s result [27] (see also Sun [37] ) to obtain the smooth solution $u^{t}$ to equation (6) with $\sup_{M}u^{t}=0$.
To prove Theorem 1.1, we establish uniform $C^{2,\alpha}$ estimates for $u^{t}$ that are independent of $t$ and $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
Theorem 1.2.
Assume the same conditions as in Theorem 1.1. Let $u^{t}$ be the smooth solution of (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $C$ and $t_{0}$, depending only on $(M,\omega,\chi)$ but independent of $t$, such that for any $t\in(0,t_{0})$,
$|u^{t}|_{C^{2,\alpha}(M)}\leq C.$ (7)
According to Collins–Jacob–Yau [7], $\sup_{M}|u^{t}|$ is uniformly bounded. Their results imply that the $C^{2,\alpha}$ estimates can be reduced to proving uniform complex Hessian estimates. Our main contribution is the proof of the uniform complex Hessian estimate of Hou–Ma–Wu type,
$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C\bigl(1+\sup\limits_{M}|\partial u|_{\omega}^{2}\bigr),$ (8)
and the uniform gradient estimate via the Dinew–Kolodziej blow-up argument [10].
However, the situation here is more subtle. If the gradient estimate fails, we obtain a bounded non-constant $C^{1}$ function $v$ on $\mathbb{C}^{n}$ that, in the sense of currents, is a weak solution to one of the following equations:
$\sigma_{n}=0,\quad\sigma_{n-1}=0,\quad\text{or}\quad\sigma_{n-1}+\sigma_{n}=0.$
The last case is new and requires a corresponding Liouville-type theorem. Note that $v+|z|^{2}$ is plurisubharmonic and $\Delta v$ is uniformly bounded in the weak sense (see Remark 4.2). When following the argument in [10, 38], we find that the first case can be handled similarly. For the second case, we need to show that the standard mollification $v]_{r}$ of $v$, as well as $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$, are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is sufficiently small. To address this issue, our starting observation is that a smooth function $v$ satisfies $\sigma_{n-1}+\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n}$ if and only if the function $v+|z_{n+1}|^{2}$ satisfies $\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n+1}$, where $z_{n+1}$ is a new complex variable. Based on this observation, we use potential theory to obtain the required Liouville-type theorem (see Theorem 4.1).
Finally, we find some applications of the main result. In general, the critical LYZ equation has the form
$\sigma_{n-1}(\chi_{u})=\sigma_{n-3}(\chi_{u})-\sigma_{n-5}(\chi_{u})+\cdots,$
where $\sigma_{k}(\chi_{u})$ is the $k$-Hessian operator of $\chi_{u}$ with respect to the metric $\omega$ (see (11)).
When $n=3$, the critical LYZ equation is just the 2-Hessian equation
$\sigma_{2}(\chi_{u})=1.$
As a consequence of our result, we can solve the 2-Hessian equation in dimension 3 under weaker conditions than those by Sun [36].
Corollary 1.1.
Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$-form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$-form and
$3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3},\ \ \int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}.$
Then there exists a unique smooth solution $u$ solving
$\sigma_{2}(\chi_{u})=1,$
$\chi_{u}\in\Gamma_{2}(M),\ \sup_{M}u=0.$
Note that Hou-Ma-Wu [18] proved the complex Hessian estimate of Hou-Ma-Wu type for the $k$-Hessian equation $\sigma_{k}(\omega_{u})=f$ on a compact Kähler manifold. Then Dinew-Kolodziej [10] obtained the gradient estimate by a Liouville theorem for $\sigma_{k}=0$ in $\mathbb{C}^{n}$ and thus solved the equation. The Hermitian case was solved independently by Székelyhidi [38] and Zhang [44].
When $n=4$, the critical LYZ equation is the Hessian quotient equation
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}).$
Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than those by Sun [36] and Székelyhidi [38].
Corollary 1.2.
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$-form}$ and
$\int_{M}\chi^{3}\wedge\omega=$ $\int_{M}\chi\wedge\omega^{3},$
$\int_{M}\chi\wedge\omega^{3}>$ $0,\ \mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0.$
Then there exists a unique smooth function $u$ solving
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}),$ (9)
$\chi_{u}\in\Gamma_{3}(M),\ \sup_{M}u=0.$
We remark that the Hessian quotient equation is of the form $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$. It reduces to the $J$-equation when $k=n$ and l=n1l=n-1 which was solved by Song-Weinkove [35] using the flow method, and the case $k=n,1\leq l\leq n-1$ was solved by Fang-Lai-Ma [11]. The general case was subsequently solved by Sun [36] and Székelyhidi [38] via different continuity methods.
The organization of this paper is as follows. In Section 2, we provide preliminaries on $k$-Hessian operators, the special Lagrangian operator $\theta_{\omega}$, and the approximating family $\theta(t)$. We also give the $C^{0}$-estimate for $u^{t}$ using results from [7]. In Section 3, we first prove a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$. In Section 4, based on the complex Hessian estimate, we prove the gradient estimate by a blow up argument. Then we prove the main theorems. In Section 5, we apply the main results to the 3-dimensional 2-Hessian equation and 4-dimensional Hessian quotient equation.
2. preliminaries
2.1. The $k$-Hessian operators and the special Lagrangian operator
For an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
$\sigma_{k}(\lambda)=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to the Kähler form $\omega$ is defined by
$\sigma_{k}(\alpha):=C_{n}^{k}\frac{\alpha^{k}\wedge\omega^{n-k}}{\omega^{n}}.$ (11)
In local coordinates, writing $\alpha=\sqrt{-1}\alpha_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$, we have
$\sigma_{k}(\alpha)=\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr).$
For simplicity, we also denote $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ for $\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr)$.
We need the following convex cones:
$\Gamma_{k}$ $:=\{\lambda\in\mathbb{R}^{n}\mid\sigma_{m}(\lambda)>0,\ 1\leq m\leq k\},$
[…]
$\mathcal{M}(\Gamma_{n-1})$ $:=\{A\mid\text{$A$ is an $n\times n$ Hermitian matrix and }\lambda(A)\in\Gamma_{n-1}\}.$
The Hessian operators possess the following concavity property.
Lemma 2.1.
The function $\sigma_{n-1}^{\frac{1}{n-1}}(\lambda)$ is concave in $\Gamma_{n-1}$. Consequently, the function $\sigma^{\frac{1}{n-1}}_{n-1}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
Let
$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\text{for }\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n},$ (12)
and define the cones
$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\text{for }\tau\in\bigl[(n-2)\tfrac{\pi}{2},n\tfrac{\pi}{2}\bigr).$
The following lemma collects some key properties of these cones.
Lemma 2.2 (Yuan [43], Wang-Yuan [41]).
Assume that $\theta(\lambda)\geq\tau$ for some $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$, and that the components of $\lambda=(\lambda_{1},\dots,\lambda_{n})$ satisfy $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$. Then the following hold:
  1. (1)
    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$ and $\lambda_{n-1}\geq|\lambda_{n}|$;
  2. (2)
    $\lambda_{1}+(n-1)\lambda_{n}\geq 0$;
  3. (3)
    $\lambda\in\Gamma_{n-1}$;
  4. (4)
    $\Gamma^{\tau}$ is convex.
$2.$ $C^{0}$-estimates
Let $\hat{\theta}(t)$ denote the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$, and set $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$. We first show that $\hat{\theta}(t)$ is close to $\pi$ when $t>0$ is sufficiently small.
Lemma 2.3.
Assume that $\underline{u}$ is a subsolution of the critical LYZ equation (4). Then there exist uniform positive constants $t_{0}$ and $C$, depending only on $\chi$ and $\omega$, such that for all $t\in(0,t_{0})$,
$\hat{\theta}(t)\in\bigl(\pi-2Ct,\;\pi-Ct\bigr).$
Proof.
By a direct computation, we have
$(\chi+t\omega+\sqrt{-1}\omega)^{n}=(\chi+\sqrt{-1}\omega)^{n}+t\,n(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2}).$
Since $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0$, it follows that for sufficiently small $t$,
$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$
We now show that the imaginary part is positive. The subsolution condition implies the positivity
$\mathrm{Im}\,(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}>0.$
Hence, for sufficiently small $t>0$,
$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0.$ (13)
Consequently,
$\mathrm{Im}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ $=\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n}$
$\quad+t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
$=t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
$>0,$
where the last inequality uses (13).
Therefore, there exists a uniform constant $C>0$, depending only on $\chi$ and $\omega$, such that the principal argument $\hat{\theta}(t)$ of $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
$\hat{\theta}(t)=\pi-\arctan\!\left(\frac{t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}{-\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}\right)\in(\pi-2Ct,\;\pi-Ct)$
for all sufficiently small $t>0$. ∎
Hence, for sufficiently small $t>0$, Lemma 2.3 yields
$\theta(t)\in\bigl((n-2)\tfrac{\pi}{2}+Ct,\;(n-2)\tfrac{\pi}{2}+2Ct\bigr),$ (14)
and consequently,
$\min_{1\leq j\leq n}\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\arctan\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2}.$ (15)
This shows that $\underline{u}$ remains a subsolution of equation (6) for all sufficiently small $t>0$.
By the result of Lin [27] (see also Sun [37]), for each such $t$, there exists a smooth solution $u^{t}$ to the supercritical LYZ equation (6). Moreover, Proposition 3.6 of [7] provides a uniform $C^{0}$ estimate for these solutions.
Lemma 2.4 (Collins–Jacob–Yau [7]).
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4), and let $u^{t}$ be the solution of (6) normalized by $\sup_{M}u^{t}=0$. Then there exist constants $t_{0}>0$ and $C>0$, depending only on the background data $(\chi,\omega)$ and $\underline{u}$, such that for all $t\in(0,t_{0})$,
$\|u^{t}\|_{C^{0}(M)}\leq C.$
3. Complex Hessian estimates
3.1. The linearized operator
We first collect some useful formulas for the complex Hessian estimate from [12].
Lemma 3.1.
Let $u^{t}$ be a solution of the LYZ equation (6). Then:
  1. (i)
    The linearized operator $\mathcal{L}$ of (6) takes the form
    $\mathcal{L}(v)=\sum_{i,j}F^{i\bar{j}}v_{i\bar{j}},$
    where the matrix $(F^{i\bar{j}})$ is the inverse of $wg^{-1}w+g$. Here $g=(g_{i\bar{j}})$ and $w=(w_{i\bar{j}})$ with $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$.
  2. (ii)
    Differentiating the LYZ equation (6) (using covariant derivatives) yields
    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=\sum_{i,j,k,l}F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}\bigl(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}\bigr).$ (17)
3.2. Key inequalities from the subsolution condition
The subsolution condition plays a crucial role in the $C^{2}$-estimates for solutions of fully nonlinear elliptic PDEs, as demonstrated in the works of Guan [13] and Székelyhidi [38]. Following the approach in these references, we establish key inequalities that will be essential for the second-order estimate. A favorable feature is that the constant $\delta_{0}$ appearing below can be chosen uniformly.
Lemma 3.2.
Let $\mu\in\mathbb{R}^{n}$ satisfy
$A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}\;>\;A_{0}\;>\;(n-3)\frac{\pi}{2}.$
Then, for every $\lambda\in\mathbb{R}^{n}$ with $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$, depending only on $n$, $\mu$, $A_{0}$, and the constant $C$ from Lemma 2.3, such that for any $t\in(0,t_{0})$, one of the following statements holds:
  1. (i)
    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$; or
  2. (ii)
    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$ for $1\leq j\leq n$.
Proof.
Given that $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta>0$ such that
$A(\mu-2\delta\mathbf{1})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
where $\mathbf{1}=(1,\ldots,1)\in\mathbb{R}^{n}$.
For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, assume without loss of generality that $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Recall that the vector
$\mathbf{n}=\Bigl(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}}\Bigr)$
is an inward normal to $\Gamma^{\theta(t)}$ at $\lambda$. Let $T_{\lambda}$ denote the tangent plane to the hypersurface $\partial\Gamma^{\theta(t)}$ at $\lambda$.
Case 1: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}\geq 0$. In this case, we have
$\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (18)
Case 2: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}<0$. We claim that there exists a positive constant $s$ such that $\mu-\delta\mathbf{1}+se_{1}\in\Gamma^{\theta(t)}$. Indeed, observe that
$\theta(\mu-\delta\mathbf{1}+se_{1})$ $=\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
$>\arctan(\mu_{1}-\delta+s)+A(\mu-\delta\mathbf{1})$
$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
To ensure $\theta(\mu-\delta\mathbf{1}+se_{1})>\theta(t)$, it suffices to choose $s$ satisfying
$\arctan(\mu_{1}-\delta+s)>\theta(t)-\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
Take
$s=|\delta-\mu_{1}|+\tan\Bigl(\frac{\pi}{2}-\frac{1}{4}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)\Bigr),$
and assume $0<t<t_{0}$ with $t_{0}=\frac{1}{8C}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)$. Then the desired inequality holds.
The convexity of $\Gamma^{\theta(t)}$ implies that the points $\mu-\delta\mathbf{1}$ and $\mu-\delta\mathbf{1}+s\mathbf{e}_{1}$ lie on opposite sides of the tangent plane $T_{\lambda}$. Hence, there exists $a_{0}\in(0,1)$ such that the convex combination
$a_{0}(\mu-\delta\mathbf{1}+s\mathbf{e}_{1})+(1-a_{0})(\mu-\delta\mathbf{1})=\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}$
lies on $T_{\lambda}$.
We now show that $\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\in\Gamma^{\theta(t)}$. Indeed,
$\theta\bigl(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\bigr)$
$=\arctan(\mu_{1}-\delta+a_{0}s+s)+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
$\geq\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\mathbf{1})$
$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr)$
$>\theta(t).$
Thus, we have
$0$ $\leq\bigl((\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n})-(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1})\bigr)\cdot\mathbf{n}$
$=(s\mathbf{e}_{1}-\delta\mathbf{e}_{n})\cdot\mathbf{n}$
$=\frac{s}{1+\lambda_{1}^{2}}-\frac{\delta}{1+\lambda_{n}^{2}},$
which is equivalent to
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{n}^{2}}.$ (19)
By Lemma 2.2, we have $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$. Consequently, for each $j=1,\dots,n$,
$\frac{1}{1+\lambda_{n}^{2}}\geq\frac{1}{1+\lambda_{j}^{2}}.$
Combining with (19), we obtain
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{j}^{2}},\quad\textup{for all $1\leq j\leq n$.}$
Summing over $j=1,\dots,n$ yields
$n\cdot\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}},$
and therefore
$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}}.$
Since $\lambda_{1}^{2}\geq\lambda_{j}^{2}$ for all $j$, we have
$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (20)
In view of (18) and (20), choosing $\delta_{0}=\frac{\delta}{ns}$ completes the proof. ∎
By the preceding lemma, we obtain the following key result, noting that the estimate does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
Lemma 3.3.
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and $u^{t}$ the solution of the LYZ equation (6). Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and any $z\in M$, one of the following holds at $z$:
  1. (i)
    $\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, or
  2. (ii)
    $\mathcal{F}_{\min}\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, where $\mathcal{F}_{\min}$ denotes the smallest eigenvalue of the Hermitian matrix $\bigl(F^{i\bar{k}}g_{j\bar{k}}\bigr)_{n\times n}$.
Proof.
Let $\mu=(\mu_{1},\dots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$, arranged so that $\mu_{1}\geq\cdots\geq\mu_{n}$.
Fix $z\in M$. We choose normal coordinates at $z$ such that
$g_{i\bar{j}}=\delta_{ij},\quad\chi_{t,u^{t}}=\sum_{i=1}^{n}\lambda_{i}\sqrt{-1}\,dz^{i}\wedge d\bar{z}^{i}\quad\text{with}\quad\lambda_{1}\geq\cdots\geq\lambda_{n},$
where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t}$. In these coordinates, at $z$ we have
$F^{i\bar{j}}=\frac{\delta_{ij}}{1+\lambda_{i}^{2}}.$ (21)
Let $\delta_{0}$ and $t_{0}$ be the uniform constants given by Lemma 3.2. We now prove the lemma by distinguishing two cases.
Case 1: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum_{i}F^{i\bar{i}}$.
Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, the Schur–Horn theorem [17] implies
$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})$ $=\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)+t\sum_{i}F^{i\bar{i}}$
$\geq\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)$
$=\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
$\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
Case 2: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum_{i}F^{i\bar{i}}$.
By Lemma 3.3, we obtain
$\mathcal{F}_{\min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}.$
This completes the proof. ∎
3.3. The complex Hessian estimate of Hou-Ma-Wu type
We use the auxiliary function introduced by Hou-Ma-Wu [18], who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifolds, to derive the complex Hessian estimates for families of LYZ equations. Here, instead of using the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$, we employ the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$, which simplifies the proof considerably.
Theorem 3.1.
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and let $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exist uniform positive constants $t_{0}$ and $C$, independent of $\bigl(\theta(t)-(n-2)\frac{\pi}{2}\bigr)^{-1}$, such that for any $t\in(0,t_{0})$,
$\sup_{M}|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}\leq C\bigl(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}\bigr).$ (22)
Proof.
Without loss of generality, we assume $\underline{u}=0$. Let $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$. For simplicity, we write $u$ instead of $u^{t}$ in the following proof. Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\mid|\xi|_{\omega}=1\}$. We consider the auxiliary function on $S(T^{1,0}M)$:
$H(z,\xi)=\log\bigl(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l}\bigr)+\varphi\bigl(|\nabla u|_{g}^{2}\bigr)+\psi(u),$
where $\varphi$ and $\psi$ are defined by
$\varphi(s)$ $=-\frac{1}{2}\log\!\left(1-\frac{s}{2K}\right),\quad 0\leq s\leq K-1,$
$\psi(s)$ $=-A\log\!\left(1+\frac{s}{2L}\right),\quad-L+1\leq s\leq 0,$
with
$K=\sup_{M}|\nabla u|_{g}^{2}+1,\quad L=\sup_{M}|u|+1,\quad A=2L(C_{0}+1),$
and $C_{0}$ is a uniform positive constant to be determined later. We have the following derivatives:
$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\quad\varphi^{\prime\prime}=2(\varphi^{\prime})^{2},$ (23)
$\frac{1}{2}AL^{-1}$ $\leq-\psi^{\prime}\leq AL^{-1},\quad\psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (24)
Suppose $H$ attains its maximum at $(z_{0},\xi_{0})$. We choose holomorphic coordinates near $z_{0}$ such that
$g_{i\bar{j}}(z_{0})=\delta_{ij},\quad\partial_{k}g_{i\bar{j}}(z_{0})=0,\quad\text{and}$
$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{ij}\quad\text{with}\quad\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n},$
which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$. Extend $\xi_{0}$ near $z_{0}$ by $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$. Then the function
$Q(z)=H(z,\tilde{\xi}_{0}(z))$
attains a local maximum at $z_{0}$.
By the maximum principle, at $z_{0}$ we have
$Q_{i}=\frac{w_{1\bar{1},i}}{w_{1\bar{1}}}+\varphi_{i}+\psi_{i}=0,$ (25)
and
$0\geq\mathcal{L}(Q)=\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}=$ $\sum_{i}\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
$+\sum_{i}\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}\sum_{i}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
$+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
We will estimate each term in (3.3). In what follows, we always assume
$\lambda_{1}>80A\sqrt{A}K,$ (27)
otherwise the desired estimate holds trivially.
Since $d\chi=0$, the commutation formula for covariant derivatives yields
$w_{1\bar{1},i\bar{i}}=w_{i\bar{i},1\bar{1}}+(\lambda_{1}-\lambda_{i})R_{1\bar{1}i\bar{i}}.$ (28)
By (17) and the fact that $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ (Lemma 2.2), we obtain
$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}$ $=\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$
$\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}+\sum_{i=1}^{n}\frac{2\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2},$ (29)
where we have used $w_{i\bar{1},1}=w_{1\bar{1},i}$, which follows from $d\chi=0$.
We now prove the following inequality
$\sum_{i}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
This is clearly true if $\lambda_{n}\geq 0$. Assume now that $\lambda_{n}<0$. By (16), we have
$\Bigl|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\Bigr|=\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|.$
Using this relation and the fact that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we obtain
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}$ $=\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$\geq\Bigl(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\Bigr)^{-1}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$=\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$\geq 0,$
where the last inequality follows from
$1+\lambda_{n}\sum_{i=1}^{n-1}\lambda_{i}^{-1}=\lambda_{n}\,\frac{\sigma_{n-1}(\lambda)}{\sigma_{n}(\lambda)}\geq 0.$
Combining (3.3) with (30), we obtain
$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
Consequently,
$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$
$\geq$ $-\lambda_{1}^{-2}(1+\lambda_{1}^{2})^{-1}|w_{1\bar{1},1}|^{2}-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
$+\sum_{i=2}^{n}\frac{\lambda_{i}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}$
$=$ $\mathcal{I}+\mathcal{II}+\mathcal{III}.$ (32)
We first estimate $\mathcal{I}$ and $\mathcal{II}$ using (25):
$\mathcal{I}$ $\geq-2(\varphi^{\prime})^{2}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}F^{1\bar{1}}|u_{1}|^{2}$
$\geq-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}},$ (33)
and
$\mathcal{II}$ $=-(1+\lambda_{1}^{2})^{-1}\sum_{i=2}^{n}\frac{F^{i\bar{i}}|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}}$
$\geq-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}$
$\quad-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
$\geq-\frac{1}{12}\varphi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-\frac{1}{2}\psi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2},$ (34)
where the last inequality follows from (27).
Next we estimate the term $\mathcal{III}$. Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we obtain
$\mathcal{III}\geq\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})}.$
Then, using (25) and the Cauchy-Schwarz inequality, we have
$\mathcal{III}$ $\geq-\frac{4}{3}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\varphi^{\prime})^{2}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
$\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2},$ (35)
where we used (23).
Let $\tau=\frac{1}{8A}$. We proceed to estimate $\mathcal{III}$ in two cases.
Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we obtain
$\mathcal{III}$ $\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-4\tau(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
$\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{2}\psi^{\prime\prime}F^{n\bar{n}}|u_{n}|^{2},$ (36)
where we used (24).
Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$, and consequently
$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq 4AF^{n\bar{n}}K\leq\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
Inserting this inequality into (3.3) yields
$\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (37)
Combining (3.3) and (37), we obtain
$\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{2}\psi^{\prime\prime}F^{n\bar{n}}|u_{n}|^{2}-\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (38)
Inserting (3.3), (3.3) and (38) into (32), we obtain
$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$ (39)
$+\frac{3}{4}\varphi^{\prime\prime}\sum_{i}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}$
$\geq$ $\frac{1}{3}\varphi^{\prime}F^{1\bar{1}}u_{1\bar{1}}^{2}-2K(\psi^{\prime})^{2}F^{1\bar{1}}>0,$ (40)
where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$, which follows from (27).
Next we estimate $\sum\limits_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=\sum_{i}F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (41)
Using the covariant derivative formulae and (16), we obtain
$\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
$\geq$ $-\sum_{i,k}F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|\text{Rm}||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
$\geq$ $-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (42)
Inserting the above estimate into (41), we get
$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (43)
Inserting (28), (39) and (43) into (3.3), we obtain
$0\geq\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}$ $\geq\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
$\geq\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (44)
By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds:
$(i)\ -\sum_{i}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i}F^{i\bar{i}},\quad\text{or}$
$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|\text{Rm}|)$.
If (i) holds, then by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we obtain a contradiction. Thus (ii) holds, and we get $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$. Consequently, we have
$0$ $\geq\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
$\geq\Bigl(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|\text{Rm}|)\Bigr)F^{n\bar{n}}.$
Therefore, $\lambda_{1}\leq CK$, where $C$ is a uniform positive constant independent of $t$. ∎
4. Gradient estimates and the proof of the main theorems
In this section, following the argument by Dinew-Kolodziej [10] and Szekelyhidi [38], we prove a Liouville theorem for the equation $\sigma_{n}+\sigma_{n-1}=0$. Then based on the uniform complex Hessian estimates of Hou-Ma-Wu type in Section 3, we get the gradient estimate by a blow-up argument. At last, we prove the main theorems.
4.1. Viscosity solutions
We begin by defining viscosity solutions for the equation
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0.$ (45)
A $C^{2}$ function $v:\Omega\rightarrow\mathbb{R}$ is said to be (n1)(n-1)-admissible in a domain $\Omega\subset\mathbb{C}^{n}$ if, for all $z\in\Omega$, the eigenvalue vector $\lambda(\sqrt{-1}\partial\bar{\partial}v(z))$ belongs to $\overline{\Gamma}_{n-1}$.
Definition 4.1.
$(i)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity subsolution of (45) if for every $C^{2}$ function $\varphi$ and every local minimum point $z_{0}\in\Omega$ of $\varphi-v$, the following inequality holds:
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))\geq 0.$
$(ii)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity supersolution of (45) if for every (n1)(n-1)-admissible function $\varphi$ and every local maximum point $z_{0}\in\Omega$ of $\varphi-v$, the following inequality holds:
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))\leq 0.$
$(iii)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity solution of (45) if it is both a viscosity subsolution and a viscosity supersolution.
If $v$ is a $C^{2}$ viscosity subsolution of (45), then it satisfies $\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)\geq 0$ in the classical sense.
We now establish some basic properties.
Lemma 4.1.
$(i)$ Let $\{v^{i}\}$ be a sequence of smooth viscosity subsolutions of (45) in $\Omega$ that converges locally uniformly to $v$. Then $v$ is also a viscosity subsolution of (45) in $\Omega$.
$(ii)$ Let $v$ be a viscosity subsolution of (45) in $\mathbb{C}^{n}$. If $v$ is independent of $z_{n}$, then its restriction to $\mathbb{C}^{n-1}$ is a viscosity subsolution of $\sigma_{n-1}=0$.
Proof.
For any $C^{2}$ function $\varphi$ and any $z_{0}\in\Omega^{n}$ of the local minimum point of $\varphi-v$, we need to show
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (46)
For any $\epsilon>0$, for all sufficiently large $i$, we have $\sup_{B_{1}(z_{0})}|v^{i}-v|<\frac{\epsilon^{3}}{3}$. Then the function $\Phi=\varphi+\epsilon|z-z_{0}|^{2}-v^{i}$ satisfies $\Phi|_{\partial B_{\epsilon}(z_{0})}>\Phi(z_{0})$, and thus $\Phi$ attains its minimum on $\overline{B_{\epsilon}(z_{0})}$ at some point $z_{i,\epsilon}\in B_{\epsilon}(z_{0})$. Since $v^{i}$ is a viscosity subsolution, we have
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
Letting $\epsilon\to 0$ then implies (46). This proves $(i)$.
$(ii)$ follows from the definition of the viscosity subsolution.
We will need the following comparison principle in the proof of the Liouville Theorem 4.1.
Lemma 4.2.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain. Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (45) in $\Omega$, and $v$ is a viscosity supersolution of (45) in $\Omega$. If $w\leq v$ on $\partial\Omega$, then $w\leq v$ in $\Omega$.
Proof.
Assume that there exists a point $x_{0}\in\Omega$ such that $w(x_{0})>v(x_{0})$. Choose $\epsilon>0$ sufficiently small so that
$\epsilon<\frac{w(x_{0})-v(x_{0})}{\max_{z\in\overline{\Omega}}|z|^{2}}.$
Since $w\leq v$ on $\partial\Omega$, we deduce the following chain of inequalities:
$\max_{z\in\overline{\Omega}}\,(w+\epsilon|z|^{2}-v)$ $\geq w(x_{0})-v(x_{0})$
$>\epsilon\cdot\max_{z\in\partial\Omega}|z|^{2}$
$\geq\max_{z\in\partial\Omega}\,(w+\epsilon|z|^{2}-v).$
This implies that the function $w+\epsilon|z|^{2}-v$ attains its maximum at an interior point $z_{0}\in\Omega$.
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity supersolution, by the definition of the supersolution, we have
$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (47)
However, since $\sigma_{k}(w_{\alpha\bar{\beta}}(z_{0}))\geq 0$ for $1\leq k\leq n-1$, expanding the left-hand side of (47) gives:
$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n}<0.$ (48)
This contradicts the assumption that $w$ is a viscosity subsolution of (45), thereby completing the proof. ∎
4.2. Weak solutions in the sense of currents
Let $\mathcal{E}=\sum_{\alpha}\sqrt{-1}dz_{\alpha}\wedge d\bar{z}_{\alpha}$ denote the standard Kähler metric on $\mathbb{C}^{n}$. We recall the definition of $k$-subharmonic functions following Blocki [1].
Definition 4.2.
For $1\leq k\leq n$, a continuous function $v:\Omega\rightarrow\mathbb{R}$ is called $k$-subharmonic if $v$ is subharmonic and
$\sqrt{-1}\partial\bar{\partial}v\wedge\gamma_{1}\wedge\cdots\wedge\gamma_{k-1}\wedge\mathcal{E}^{n-k}\geq 0\quad\text{in }\Omega$
for all real $(1,1)$-forms $\gamma_{1},\dots,\gamma_{k-1}$ with $\lambda(\gamma_{i})\in\overline{\Gamma}_{k}$ for $1\leq i\leq k-1$.
If $v$ is $C^{2}$ and $k$-subharmonic, then $v$ is $k$-admissible. More generally, by Lemma 3.7 in [30], $v$ is a viscosity subsolution of $\sigma_{k}=0$ if and only if $v$ is $k$-subharmonic.
The following lemma, which generalizes the Garding inequality, is particularly useful for $k$-subharmonic functions (see [1, page 1444, (3.4)]).
Lemma 4.3.
Let $v,v_{1},\dots,v_{k-1}$ be continuous $k$-subharmonic functions. Then
$(\sqrt{-1}\partial\bar{\partial}v)\wedge(\sqrt{-1}\partial\bar{\partial}v_{1})\wedge\cdots\wedge(\sqrt{-1}\partial\bar{\partial}v_{k-1})\wedge\mathcal{E}^{n-k}\geq 0$
in the sense of currents. In particular,
$\sqrt{-1}\partial v\wedge\bar{\partial}v\wedge(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge\mathcal{E}^{n-k}\geq 0$
in the sense of currents.
We now define weak solutions in the sense of currents.
Definition 4.3.
Let $v$ be a continuous (n1)(n-1)-subharmonic function in $\mathbb{C}^{n}$ such that $v+|z|^{2}$ is plurisubharmonic. We say that $v$ is a weak solution (respectively, weak subsolution) of the equation $\sigma_{n}+\sigma_{n-1}=0$ if it satisfies
[…]
in the sense of currents.
Remark 4.1.
The condition that $v+|z|^{2}$ is plurisubharmonic ensures that $(\sqrt{-1}\partial\bar{\partial}v)^{n}$ is well-defined as a current. Note that the blow-up solution constructed in Subsection 4.4 satisfies this condition. On the other hand, if $v$ is of class $C^{2}$, then $\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)=0$ (respectively, $\geq 0$) is equivalent to $(\sqrt{-1}\partial\bar{\partial}v)^{n}+n(\sqrt{-1}\partial\bar{\partial}v)^{n-1}\wedge\mathcal{E}=0$ (respectively, $\geq 0$).
Let $\eta\in C^{\infty}(\mathbb{C}^{n},\mathbb{R})$ be a mollifier, defined by
$\eta(z)=\begin{cases}C_{n}e^{\frac{1}{|z|^{2}-1}},&|z|<1,\\ 0,&|z|\geq 1,\end{cases}$ (49)
where the constant $C_{n}$ is chosen so that $\int_{\mathbb{C}^{n}}\eta(z)\,\mathcal{E}^{n}(z)=1$. Similarly, let $\eta_{0}$ be a mollifier on $\mathbb{C}$.
For a function $w:\mathbb{C}^{n}\rightarrow\mathbb{R}$, we define its mollification by
$[w]_{r}(z)=\int_{\mathbb{C}^{n}}w(z+ry)\eta(y)\,\mathcal{E}^{n}(y).$ (50)
For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define
[…]
where $\tilde{\eta}(\tilde{y})=\eta_{0}(y_{n+1})\eta(y)$ with $\tilde{y}=(y,y_{n+1})$. A direct computation shows that
$\langle\tilde{w}\rangle_{r}(\tilde{z})=[w]_{r}(z)+|z_{n+1}|^{2}+r^{2}M,$
where $M:=\int_{\mathbb{C}}|y_{n+1}|^{2}\eta_{0}(y_{n+1})\,\sqrt{-1}dy_{n+1}\wedge d\bar{y}_{n+1}$.
The following lemma establishes key properties of weak solutions to the equation $\sigma_{n}+\sigma_{n-1}=0$.
Lemma 4.4.
Let $v$ be a weak solution (in the sense of currents) to $\sigma_{n}+\sigma_{n-1}=0$. Then the following properties hold:
  1. (i)
    The functions $\tilde{v}(\tilde{z})=v(z)+|z_{n+1}|^{2}$ is a weak solution to $\sigma_{n}=0$ and $\langle\tilde{v}\rangle_{r}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$.
  2. (ii)
    The mollification $v]_{r}$ is a smooth subsolution to the same equation.
  3. (iii)
    If $0\leq v\leq 1$, $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
  4. (iv)
    If $w$ is a subsolution to the same equation, then $\frac{1}{2}(v+w)$ is also a subsolution.
Proof.
We first note that $v]_{r}$ is (n1)(n-1)-subharmonic since $v$ is (n1)(n-1)-subharmonic. To prove $(i)$, we will show that $\tilde{v}$ satisfies the equation: $\sigma_{n}=0$ on $\mathbb{C}^{n+1}$ in the current sense.
Let $e_{n+1}=\sqrt{-1}dy_{n+1}\wedge d\bar{y}_{n+1}$ and let $\widetilde{\mathcal{E}}=e_{n+1}+\mathcal{E}$ be the standard Kähler metric on $\mathbb{C}^{n+1}$. Since $v+|z|^{2}$ is plurisubharmonic in $\mathbb{C}^{n}$, it follows that $\tilde{v}+|z|^{2}$ is also plurisubharmonic in $\mathbb{C}^{n+1}$. Moreover, for any $1\leq k\leq n$, we have:
$(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $C^{k}_{n+1}(\sqrt{-1}\partial\bar{\partial}v+e_{n+1})^{k}\wedge(\mathcal{E}+e_{n+1})^{n+1-k}$
$=$ $\left((\sqrt{-1}\partial\bar{\partial}v)^{k}+k(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge e_{n+1}\right)\wedge\left(\mathcal{E}^{n-k+1}+(n+1-k)\mathcal{E}^{n-k}\wedge e_{n+1}\right)$
$=$ $\left((n-k+1)(\sqrt{-1}\partial\bar{\partial}v)^{k}\wedge\mathcal{E}^{n-k}+k(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge\mathcal{E}^{n-k+1}\right)\wedge e_{n+1}.$
[…]
$(\sqrt{-1}\partial\bar{\partial}v)^{n}+n(\sqrt{-1}\partial\bar{\partial}v)^{n-1}\wedge\mathcal{E}=0.$
This proves that $\tilde{v}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$ and satisfies $\sigma_{n}=0$ in the current sense. Consequently, $\langle\tilde{v}\rangle_{r}$ is also $n$-subharmonic in $\mathbb{C}^{n+1}$.
Since $\langle\tilde{v}\rangle_{r}=[v]_{r}(z)+|z_{n+1}|^{2}+r^{2}M$, it follows that $v]_{r}$ is a subsolution to the equation $\sigma_{n}+\sigma_{n-1}=0$. This establishes $(ii)$.
To prove (iv), let $W=\frac{v+w}{2}$. Then $W+|z|^{2}$ is plurisubharmonic and (n1)(n-1)-subharmonic. Define $\widetilde{W}(\tilde{z})=W(z)+|z_{n+1}|^{2}$. By $(i)$, we have $\widetilde{W}=\frac{\tilde{v}+\tilde{w}}{2}$, which is $n$-subharmonic in $\mathbb{C}^{n+1}$. Therefore, $W$ is a subsolution, proving $(iv)$.
We now prove $(iii)$ using $(i)$. Define $\tilde{F}(\tilde{z})=\frac{1}{2}v(z)^{2}+|z_{n+1}|^{2}$. Then for any $1\leq k\leq n$, we have:
$(\sqrt{-1}\partial\bar{\partial}\tilde{F})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $\left(v\sqrt{-1}\partial\bar{\partial}\tilde{v}+(1-v)e_{n+1}+\sqrt{-1}\partial v\wedge\bar{\partial}v\right)^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $v^{k}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$+kv^{k-1}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k-1}\wedge\left((1-v)e_{n+1}+\sqrt{-1}\partial v\wedge\bar{\partial}v\right)\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$+k(k-1)(1-v)v^{k-2}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k-2}\wedge e_{n+1}\wedge(\sqrt{-1}\partial v\wedge\bar{\partial}v)\wedge\widetilde{\mathcal{E}}^{n+1-k}.$
This expression is nonnegative by Lemma 4.3, since $\tilde{v}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$ and $0\leq v\leq 1$. ∎
We immediately conclude that a weak solution (subsolution) in the sense of currents is also a viscosity solution (subsolution).
Lemma 4.5.
Let $v$ be a weak subsolution in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity subsolution.
Proof.
By part $(i)$ of Lemma 4.4, we know that $\tilde{v}=v+|z_{n+1}|^{2}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$, and hence $\tilde{v}$ is a viscosity subsolution to $\sigma_{n}=0$ in $\mathbb{C}^{n+1}$.
Now, let $\varphi$ be any $C^{2}$ function and let $z_{0}$ be a local minimum point of $\varphi-v$. Consider the function $\tilde{\varphi}=\varphi+|z_{n+1}|^{2}$. Then $(z_{0},0)$ is a local minimum point of $\tilde{\varphi}-\tilde{v}$. Since $\tilde{v}$ is a viscosity subsolution to $\sigma_{n}=0$, we have
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(z_{0},0)\geq 0.$
Observing that
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(z_{0},0)=\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi)(z_{0})+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi)(z_{0}),$
we conclude that $v$ is a viscosity subsolution.
4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
In this subsection, based on the lemmas established in the preceding subsections and following arguments similar to those in [10, 38], we prove the following Liouville-type theorem.
Theorem 4.1.
Let $v:\mathbb{C}^{n}\rightarrow\mathbb{R}$ be a $C^{1}$ function that is (n1)(n-1)-subharmonic and such that $v+|z|^{2}$ is plurisubharmonic. Assume there exists a positive constant $C$ such that $\|v\|_{C^{1}(\mathbb{C}^{n})}\leq C$ and $\Delta v\leq C$ in the weak sense (see Remark 4.2 for details). If $v$ is a weak solution in the sense of currents and a viscosity supersolution to the equation
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0,$ (51)
then $v$ must be a constant.
Remark 4.2.
The condition $\Delta v\leq C$ in the weak sense means that there exists a constant $C>0$ such that for every nonnegative test function $f\in C_{c}^{\infty}(\mathbb{C}^{n})$ with $\int_{\mathbb{C}^{n}}f(z)\,\mathcal{E}^{n}(z)=1$, we have
$\int_{\mathbb{C}^{n}}v\,\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}\leq C.$
From this condition, it follows immediately that the regularization $v]_{r}$ has bounded Laplacian and hence bounded complex Hessian.
Proof of Theorem 4.1.
Assume that $v$ is not constant. Then $\inf_{\mathbb{C}^{n}}v<\sup_{\mathbb{C}^{n}}v$. By scaling, we may assume that $v$ satisfies the following normalized conditions:
$\begin{cases}\sup_{\mathbb{C}^{n}}v=1,\quad\inf_{\mathbb{C}^{n}}v=0,\\ |\nabla v|\leq C_{1}:=(\sup_{\mathbb{C}^{n}}v-\inf_{\mathbb{C}^{n}}v)^{\frac{1}{2}},\quad\Delta v\leq C,\\ \sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0.\end{cases}$ (52)
Indeed, the function $\tilde{v}(z)=C_{1}^{-2}({v(C_{1}z)-\inf_{\mathbb{C}^{n}}v})$ satisfies these conditions.
For sufficiently small $\epsilon>0$, let $v^{\epsilon}=[v]_{\epsilon}$. Consider the function
Case 1: There exist a positive constant $\rho$, sequences $\epsilon_{k}\to 0$, $x_{k}\in\mathbb{C}^{n}$, $r_{k}\to\infty$, and a sequence of unit $(1,0)$-vectors $\xi_{k}$ such that
$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(x_{k})+[v]_{\rho}(x_{k})-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (53)
$\lim_{k\to\infty}\int_{B_{r_{k}}(x_{k})}|v_{\xi_{k}}^{\epsilon}|^{2}(z)\,\mathcal{E}^{n}(z)$ $=0.$ (54)
Following the arguments in [38] and [10], we obtain an (n1)(n-1)-subharmonic function $v^{\infty}$ that is also a solution to the same equation and is independent of $z_{n}$. This implies that $v^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n-1}$. Since $v^{\infty}$ is bounded, it must be constant, contradicting (53).
Case 2: Suppose Case 1 does not hold. Then for every $\rho>0$, there exists a constant $C_{\rho}>0$ with the following property: for all $\epsilon<C_{\rho}^{-1}$, $r>C_{\rho}$, $x\in\mathbb{C}^{n}$, and any unit $(1,0)$-vector $\xi$, if $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)-v(x)\geq 0$, then
$\int_{B_{r}(x)}|v^{\epsilon}_{\xi}|^{2}(z)\,\mathcal{E}^{n}(z)\geq c_{0}:=C_{\rho}^{-1}.$ (55)
Define the open set
$\Omega:=\bigl\{z\in\mathbb{C}^{n}:v(z)<\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)-\frac{7}{6}\bigr)\bigr\},$
where $\tau_{0}>0$ is a constant to be determined later.
Since $0\leq v\leq 1$, the set $\Omega$ is bounded. As $\inf_{\mathbb{C}^{n}}v=0$, we may assume (after translation) that $v(0)<\frac{1}{60}$. By Cartan’s lemma (see [10]), we choose $\rho$ large enough so that $v]_{\rho}(0)>\frac{9}{10}$, and choose $r$ large enough so that $\bigl[\frac{v^{2}}{3}\bigr]_{r}(0)>\frac{3}{10}$. Then $0\in\Omega$.
We now show that $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)$ is a subsolution to $\sigma_{n-1}+\sigma_{n}=0$. Assuming this, the comparison principle (Lemma 4.2) implies that $v\geq w$ in $\Omega$, contradicting the definition of $\Omega$.
By Lemma 4.4 $(ii)$, we know that $v]_{r}$, $v]_{\rho}$, $\frac{(v^{\epsilon})^{2}}{2}$, and $\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}$ are all subsolutions to (51).
From (55) and following [10, 38], there exists a positive constant $b_{0}$, depending only on $c_{0}$ and $C_{n}$, such that $\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}-b_{0}|z|^{2}$ is (n1)(n-1)-subharmonic. By the concavity of $\sigma_{n-1}^{1/(n-1)}$ in $\Gamma_{n-1}$, we have
$\sigma_{n-1}^{1/(n-1)}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}\bigr)\geq(n-1)^{1/(n-1)}b_{0},$
and thus
$\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}\bigr)\geq(n-1)b_{0}^{n-1}.$ (56)
We now show that $F:=\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}-\tau_{0}|z|^{2}$ is a subsolution to $\sigma_{n}+\sigma_{n-1}=0$ for sufficiently small $\tau_{0}$. Indeed, by (56), we have:
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}F)+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}F)$
$=\sigma_{n}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+O(\tau_{0})$
$=\bigl(\frac{2}{3}\bigr)^{n}\bigl[\sigma_{n}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+\frac{3}{2}\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)\bigr]+O(\tau_{0})$
$\geq\bigl(\frac{2}{3}\bigr)^{n-1}\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{v^{2}}{3}\bigr]_{r}\bigr)+O(\tau_{0})$
$\geq\bigl(\frac{2}{3}\bigr)^{n-1}(n-1)b_{0}^{n-1}-C\tau_{0}>0,$
provided $\tau_{0}<C^{-1}\left(\frac{2b_{0}}{3}\right)^{n-1}$. Here we have used the fact that $(v^{\epsilon})^{2}]_{r}$ has uniform $C^{1}$-bounds and complex Hessian bounds.
By Lemma 4.4 $(iv)$, we get $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)$ is a subsolution to the equation $\sigma_{n}+\sigma_{n-1}=0$. ∎
4.4. Gradient estimates
We prove the gradient estimate by contradiction, following the arguments in [10, 38]. Assume that there exist a sequence $t_{i}\to 0$ and corresponding solutions $u^{t_{i}}$ to the LYZ equation (6) such that for points $p_{i}\in M$, we have
$\sup_{M}|\nabla u^{t_{i}}|_{\omega}=|\nabla u^{t_{i}}|_{\omega}(p_{i})=M_{i}\to\infty\quad\text{as }i\to\infty.$ (57)
Since $M$ is compact, by passing to a subsequence, we may assume that $\{p_{i}\}$ converges to a point $p_{0}\in M$. We choose local coordinates $\{w_{\alpha}\}_{1\leq\alpha\leq n}$ in a ball $B_{r}(p_{0})$ such that $p_{i}\in B_{r/2}(p_{0})$ and $\omega(p_{0})=\sum_{\alpha}\sqrt{-1}dw_{\alpha}\wedge d\bar{w}_{\alpha}$.
For any $R>0$, we can choose $i$ sufficiently large such that for all $z\in B_{R}(0)$, the point $p_{i}+\frac{z}{M_{i}}$ lies in $B_{2r/3}(p_{0})$. We define the rescaled functions
$\hat{u}^{t_{i}}(z)=u^{t_{i}}\bigl(p_{i}+\frac{z}{M_{i}}\bigr).$
We adopt the following notation:
$\hat{u}^{t_{i}}_{\alpha}(z)=\frac{\partial\hat{u}^{t_{i}}}{\partial z_{\alpha}}(z),\quad u^{t_{i}}_{\alpha}(w)=\frac{\partial u^{t_{i}}}{\partial w_{\alpha}}(w),$
$\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z)=\frac{\partial^{2}\hat{u}^{t_{i}}}{\partial z_{\alpha}\partial\bar{z}_{\beta}}(z),\quad u^{t_{i}}_{\alpha\bar{\beta}}(w)=\frac{\partial^{2}u^{t_{i}}}{\partial w_{\alpha}\partial\bar{w}_{\beta}}(w).$
Then we have the scaling relations:
$u^{t_{i}}_{\alpha}(w)=M_{i}\hat{u}^{t_{i}}_{\alpha}(z),\quad u^{t_{i}}_{\alpha\bar{\beta}}(w)=M_{i}^{2}\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z).$ (58)
By the $C^{0}$-estimate and the complex Hessian estimate for $u^{t_{i}}$ (for large $i$), we obtain
$\sup_{B_{R}(0)}\left(|\hat{u}^{t_{i}}|+|\partial\hat{u}^{t_{i}}|_{\mathcal{E}}+|\Delta\hat{u}^{t_{i}}|_{\mathcal{E}}\right)\leq C.$ (59)
By standard elliptic theory, we have the estimates
$|\hat{u}^{t_{i}}|_{C^{1,\frac{2}{3}}(B_{2R/3}(0))}\leq C.$ (60)
Passing to a subsequence, we find that $\hat{u}^{t_{i}}\to u^{\infty}$ in $C^{1,\frac{1}{2}}(B_{R/2}(0))$. Using a diagonal subsequence argument, we conclude that $\hat{u}^{t_{i}}\to u^{\infty}$ uniformly on compact subsets of $\mathbb{C}^{n}$, and $u^{\infty}$ is a bounded $C^{1,\frac{1}{2}}$ function defined on $\mathbb{C}^{n}$.
Since $u^{t_{i}}$ solves (6), by Lemma 2.2, we have $\sigma_{k}(\chi+t_{i}\omega+\sqrt{-1}\partial\bar{\partial}u^{t_{i}})\geq 0$ for $1\leq k\leq n-1$. It follows that $u^{\infty}$ is (n1)(n-1)-subharmonic. By construction, we have $|\partial u^{\infty}(0)|=1$, so $u^{\infty}$ is nonconstant.
Now, let
$a_{0}=-\liminf_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$
Recall that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
$a_{0}=-\lim_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$ (61)
Case 1: $a_{0}=+\infty$. In this case, we show that $u^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n}$. Since $u^{t_{i}}$ solves the LYZ equation (6), we have
$\chi+t_{i}\omega+\sqrt{-1}\partial\bar{\partial}u^{t_{i}}>\cot\hat{\theta}(t_{i})\omega.$
Then in the coordinate ball $B_{r}(p_{0})$,
$(u^{t_{i}}_{\alpha\bar{\beta}})_{n\times n}\geq(\cot\hat{\theta}(t_{i})-C)g.$ (62)
Now fix $R>0$ and $\epsilon>0$. For sufficiently large $i$ and all $z\in B_{R}(0)$, we have
$(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z))_{n\times n}\geq-M_{i}^{-2}(\cot\hat{\theta}(t_{i})-C)g\geq-\epsilon I_{n},$ (63)
where we have used the fact that $\lim_{i\to\infty}M_{i}^{-2}\cot\hat{\theta}(t_{i})=0$. Therefore, $\hat{u}^{t_{i}}+\epsilon|z|^{2}$ is plurisubharmonic in $B_{R}(0)$. Since $\hat{u}^{t_{i}}$ converges uniformly to $u^{\infty}$ on compact sets, it follows that $u^{\infty}+\epsilon|z|^{2}$ is plurisubharmonic in $B_{R}(0)$.
As $R>0$ and $\epsilon>0$ are arbitrary, we conclude that $u^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n}$. However, since $u^{\infty}$ is bounded and plurisubharmonic on $\mathbb{C}^{n}$, it must be constant. This contradicts the fact that $|\partial u^{\infty}(0)|=1$.
Case 2: $0\leq a_{0}<\infty$. Define
$A^{i}(z)=\left(g^{-1}(p_{i}+M_{i}^{-1}z)-I_{n}\right)(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z))_{n\times n}+M_{i}^{-2}g^{-1}\chi(p_{i}+M_{i}^{-1}z)+t_{i}M_{i}^{-2}I_{n}.$
Then $A^{i}$ converges uniformly to 0 on any compact subset of $\mathbb{C}^{n}$.
Since $u^{t_{i}}$ solves the LYZ equation (6), we have
$\sigma_{n-1}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)=M_{i}^{2}\tan\hat{\theta}(t_{i})\sigma_{n}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)+O(M_{i}^{-2}),$ (64)
where the matrix $A^{i}$ is uniformly bounded and by Lemma 2.2, we have
$\sigma_{k}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)>0\quad\text{for any }1\leq k\leq n-1.$ (65)
Subcase 2.1: $a_{0}=0$. In this subcase, by (64), (65), and the fact that $\hat{u}^{t_{i}}$ converges locally uniformly to $u^{\infty}$, we conclude that $u^{\infty}$ is (n1)(n-1)-subharmonic and satisfies $\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}u^{\infty})=0$ in the sense of currents. Hence, applying the Liouville theorem of Dinew-Kolodziej [10], we find that $u^{\infty}$ is constant, which is a contradiction.
Subcase 2.2: $0<a_{0}<\infty$. Since $\hat{u}^{t_{i}}$ converges uniformly to $u^{\infty}$ on compact subsets of $\mathbb{C}^{n}$, we have that $u^{\infty}+a_{0}|z|^{2}$ is plurisubharmonic, (n1)(n-1)-subharmonic, and $u^{\infty}$ satisfies the equation $\sigma_{n}+a_{0}\sigma_{n-1}=0$ in the sense of currents.
Moreover, for any test function $f\in C_{c}^{\infty}(\mathbb{C}^{n})$ with $f\geq 0$ and $\int_{\mathbb{C}^{n}}f=1$, we have
$\int_{\mathbb{C}^{n}}u^{\infty}\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}$ $=\lim_{i\to\infty}\int_{\mathbb{C}^{n}}\hat{u}^{t_{i}}\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}$
$=\lim_{i\to\infty}\int_{\mathbb{C}^{n}}f\sqrt{-1}\partial\bar{\partial}\hat{u}^{t_{i}}\wedge\mathcal{E}^{n-1}$
$\leq C.$
This proves that $\Delta u^{\infty}\leq C$ in the weak sense.
Next, we prove $u^{\infty}$ is a supersolution to the equation $\sigma_{n-1}+a_{0}\sigma_{n}=0$ in the viscosity sense. For any $C^{2}$ function $\varphi$ with $\lambda(\varphi_{\alpha\bar{\beta}})\in\overline{\Gamma}_{n-1}$ and for any $z_{0}$ of the local maximum point of $\varphi-u^{\infty}$, we need to show
$\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)+a_{0}\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)\leq 0.$ (66)
If $\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))=0$, then $\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\leq 0$ and thus (66) holds.
So we assume $c_{0}=\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))>0$. Then for any $\epsilon>0$ small enough, we have
$\lambda(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})\in\Gamma_{n-1}\ \ \text{and}\ \ \sigma_{n-1}(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})>\frac{c_{0}}{2}\ \text{in}\ \overline{B}_{\epsilon}(z_{0}).$ (67)
For sufficiently large $i$ such that $|\hat{u}^{t_{i}}-u^{\infty}|<\frac{\epsilon^{3}}{2}$, consider the function
$\Phi=\varphi-\epsilon|z-z_{0}|^{2}-\hat{u}^{t_{i}}.$
Since for any $z\in\partial B_{\epsilon}(z_{0})$,
$\Phi(z)=\varphi(z)-\hat{u}^{t_{i}}(z)-\epsilon^{3}<$ $\varphi(z)-u^{\infty}(z)-\frac{\epsilon^{3}}{2}$
$\leq$ $\varphi(z_{0})-u^{\infty}(z_{0})-\frac{\epsilon^{3}}{2}<\varphi(z_{0})-\hat{u}^{t_{i}}(z_{0})=\Phi(z_{0}).$
then we know that $\Phi$ attains its maximum on $\overline{B}_{\epsilon}(z_{0})$ at an interior point $\tilde{z}_{0}\in B_{\epsilon}(z_{0})$. Then $(\Phi_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\leq 0$, namely $(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-\epsilon\delta_{\alpha\beta})_{n\times n}\leq(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}$. Thus $(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\geq(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta})_{n\times n}$ and we obtain
$\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A_{\alpha\bar{\beta}}(\tilde{z}_{0}))\geq\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)\geq\frac{c_{0}}{2}.$
By the concavity of $\frac{\sigma_{n}}{\sigma_{n-1}}$ in $\Gamma_{n-1}$, we get
$1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}{\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}$
$\leq$ $1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}{\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}$
$\leq$ $c_{0}^{-1}CM_{i}^{-2},$ (68)
where in the last inequality we use (64). Then we get
${\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}-\tan\hat{\theta}(t_{i})M^{2}_{i}{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}\leq c_{0}^{-1}CM_{i}^{-2}\sup|D^{2}\varphi|.$ (69)
Let $i\rightarrow\inftyand$ then $\epsilon\rightarrow 0$ in the above, we obtain (66). Hence $u^{\infty}$ is a supersolution to the equation $\sigma_{n}+\sigma_{n-1}=0$.
We can then apply the Liouville Theorem 4.1 to $a_{0}^{-1}u^{\infty}$ to conclude that $a_{0}^{-1}u^{\infty}$ is constant, which again leads to a contradiction.
4.5. Proof of Theorems 1.1 and 1.2
Once the gradient estimates for $u^{t}$ are established, we obtain the uniform bound
$\|u^{t}\|_{C^{1}}+\max_{M}|\partial\bar{\partial}u^{t}|\leq C,$
where $C$ is independent of $t$. Following the argument of Collins–Jacob–Yau [7], we derive uniform $C^{2,\alpha}$ estimates for $u^{t}$, also independent of $t$. For any $\beta\in(0,\alpha)$, there exists a subsequence $\{u^{t_{i}}\}$ converging in $C^{2,\beta}$ to a limit $u^{0}\in C^{2,\alpha}$. Thus, $u^{0}$ solves the critical LYZ equation (4). By standard elliptic regularity theory, $u^{0}$ is smooth.
5. Examples: the 3-dimensional case and 4-dimensional case
In this section, as corollaries of our main theorem, we solve the 2-Hessian equation in dimension 3 and a Hessian quotient equation in dimension 4 under weaker conditions than those in the previous works.
5.1. 3-dimensional case
In this case, since
$(\chi_{u}+\sqrt{-1}\omega)^{3}=$ $\chi_{u}^{3}-3\chi_{u}\wedge\omega^{2}+\sqrt{-1}(3\chi_{u}^{2}\wedge\omega-\omega^{3}),$ (70)
the critical LYZ equation is exactly the 2-Hessian equation
$3\chi_{u}^{2}\wedge\omega=\omega^{3}.$
Sun [36] solved the 2-Hessian equation under the following conditions:
$3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3},\quad\text{and}\quad\chi\in\Gamma_{2}(M),$
with the following continuity method
$3\chi_{u}^{2}\wedge\omega=\Big(t+(1-t)\frac{3\chi^{2}\wedge\omega}{\omega^{3}}\Big)\omega^{3}.$ (71)
We consider a family of LYZ equations (6)
$3(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})^{2}\wedge\omega=$ $\tan\hat{\theta}(t)(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})^{3}$
$-3\tan\hat{\theta}(t)(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})\wedge\omega^{2}+\omega^{3}.$
As a corollary of our theorem, we solve the 2-Hessian equation under weaker conditions instead of the usual condition $\chi\in\Gamma_{2}(M)$.
Corollary 5.1.
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$-form satisfying the following conditions:
  1. (i)
    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$;
  2. (ii)
    the subsolution condition
    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$-form}.$ (72)
Then there exists a unique smooth solution $u$ solving
$\sigma_{2}(\chi_{u})=1,$ (73)
$\chi_{u}\in\Gamma_{2}(M),\ \sup_{M}u=0.$ (74)
Proof.
Condition $(i)$ gives that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$. Condition $(ii)$ means that 0 is a subsolution of the critical LYZ equation in dimension $3$. According to Theorem 1.1, there exists a smooth solution to the critical LYZ equation in dimension $3$, which is exactly the $2$-Hessian equation $\sigma_{2}(\chi_{u})=1$. ∎
Remark 5.1.
If $\chi\in\Gamma_{2}(M)$, then it satisfies the subsolution condition $(ii)$. However, the subsolution condition is strictly weaker than $\chi\in\Gamma_{2}(M)$. We will prove that the integral condition $(i)$ is necessary.
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$. Since $\sigma_{2}(\chi_{u})=1$, we have
$\sigma_{3}(\chi_{u})\leq\frac{1}{9}\sigma_{1}(\chi_{u})\sigma_{2}(\chi_{u})=\frac{1}{9}\sigma_{1}(\chi_{u}).$ (75)
Integrating by parts, we obtain
$\int_{M}\chi^{3}$ $=\int_{M}\chi_{u}^{3}\leq\frac{1}{3}\int_{M}{\chi_{u}\wedge\omega^{2}}=\frac{1}{3}\int_{M}{\chi\wedge\omega^{2}}$
$<3\int_{M}{\chi\wedge\omega^{2}},$
which proves the inequality in condition $(i)$.
5.2. $4$-dimesional case
When $n=4$, the critical LYZ equation is the Hessian quotient equation. In this case, the subsolution condition yields $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form. Under the additional assumption that $\chi\in\Gamma_{3}(M)$, Sun [36] and Székelyhidi [38] proved the existence result using different continuity methods.
As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under conditions weaker than those previously required.
Corollary 5.2.
Let $(M,\omega)$ be a compact Kähler manifold of dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
  1. (i)
    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3}$, $\int_{M}\chi\wedge\omega^{3}>0$, and $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0$;
  2. (ii)
    $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form.
Then there exists a unique smooth function $u$ solving
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}),$ (76)
$\chi_{u}\in\Gamma_{3}(M),\quad\sup_{M}u=0.$
Proof.
Condition $(i)$ implies that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$. To apply Theorem 1.1, we only need to show that 0 is a subsolution of the critical LYZ equation.
By condition $(ii)$, we have $\sigma_{2}(\chi)>2$, which implies $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$. Moreover, since there exists a point $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we conclude that $\sigma_{1}(\chi)>0$ on $M$.
Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$, ordered such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
$\arctan\mu_{2}+\arctan\mu_{3}+\arctan\mu_{4}>\frac{\pi}{2}.$ (77)
Indeed, condition $(ii)$ gives $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$. Since $\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$, it follows that $\mu_{2}+\mu_{3}>0$, and consequently
$\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}.$
Therefore,
$\arctan\mu_{4}>\arctan\left(\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}\right)=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
which proves (77). Hence, 0 is a subsolution of the critical LYZ equation. ∎
We next show that condition $(i)$ is necessary.
Lemma 5.1.
Let $u$ be the unique smooth solution to the Hessian quotient equation (76). Then $\chi$ satisfies condition $(i)$.
Proof.
The equality in condition $(i)$ follows directly from integrating the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
$\int_{M}\chi\wedge\omega^{3}=\int_{M}\chi_{u}\wedge\omega^{3}>0.$ (78)
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$, ordered such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$.
By the MacLaurin inequality, we obtain
$\sigma_{2}\geq 6\frac{\sigma_{3}}{\sigma_{1}}=6.$ (79)
The inequality $\lambda_{4}\sigma_{1}(\lambda|4)\leq\sigma_{2}(\lambda|4)$ yields
$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|4)\leq 2\sigma_{2}(\lambda|4),$ (80)
and consequently,
$\sigma_{2}(\lambda|4)\geq 3.$ (81)
Using the identity $\lambda_{4}=\dfrac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we compute
$\sigma_{4}-\sigma_{2}+1$ $=\lambda_{4}\sigma_{3}(\lambda|4)-\lambda_{4}\sigma_{1}(\lambda|4)-\sigma_{2}(\lambda|4)+1$
$=\frac{\bigl(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\bigr)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
$\leq-2.$
Therefore,
$\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ $=\mathrm{Re}\int_{M}(\chi_{u}+\sqrt{-1}\omega)^{4}$
$=\int_{M}(\sigma_{4}-\sigma_{2}+1)\,\omega^{4}<0,$
which establishes the second inequality in condition $(i)$. ∎